OpenStudy (willa):

How do I figure this out???? Which of the following is equivalen to 6x2 - 11x - 2? A) (6x - 1)(x + 2) B) (3x - 1)(2x + 2) C) (3x + 1)(2x - 2) D) (6x + 1)(x - 2) E) (x + 1)(6x - 2)

OpenStudy (anonymous):

first and last products are all correct so you just have to worry about the 'middle terms' \[12x-x=11x\] no \[6x-2x=4x\] no \[-6x+2x=-4x\] nope \[-12x+x=-11x\]got it!

OpenStudy (anonymous):

yeah thats the aswer satellite73

OpenStudy (willa):

thank you

OpenStudy (anonymous):

:)

OpenStudy (willa):

How would I solve it if it wasnt multiple choice???

OpenStudy (anonymous):

grind it until you find it.

OpenStudy (willa):

ok thanks